OpenStudy (anonymous):

Find the y-intercept of the following function. f(x) = -2x2 - 4x

OpenStudy (anonymous):

0

OpenStudy (anonymous):

f(0) is the y int f(0)---plug in 0 for all x f(0)=0

OpenStudy (anonymous):

it's a downward facing parabola with its vertex at the origin

OpenStudy (anonymous):

I told you arricuhh :P <3 set x = 0 and solve for y f(x) = -2x^2 - 4x f(0) = -2(0)^2 - 4(0) f(0) = 0 y-int = 0 :D